OpenStudy (anonymous):

Solve for x: −4|x + 5| = −16 x = 21 over 4, x = −11 over 4 x = −1, x = 9 x = −1, x = −9 No solution

OpenStudy (haseeb96):

no solution ......is the correct answer

OpenStudy (goldphenoix):

It is against the CoC to give out direct answer. Please don't do it ever again.

OpenStudy (phi):

First, simplify −4|x + 5| = −16 divide both sides by -4 what do you get ?

OpenStudy (anonymous):

-4 divide by negitive 4 is 1 and 16 divided by -4 is 4

OpenStudy (phi):

yes, you get 1*|x+5|= 4 but 1 times anything is the "anything", so that is just |x+5|= 4 now you need to "get rid of " the absolute value signs.

OpenStudy (phi):

to get rid of the | | we create TWO different equations equation one: drop the | | (that is easy) you get x+5= 4 equation 2: change the | | into parens, and multiply by -1 -1*(x+5) = 4 can you solve either of these equations ?

OpenStudy (anonymous):

i dont think so

OpenStudy (phi):

let's try x+5=4 add -5 to both sides what do you get ?

OpenStudy (anonymous):

-1

OpenStudy (phi):

you should write out the whole equation x= -1 that is one of the answers. now let's try -1*(x+5) = 4 multiply both sides by -1 what do you get (write out the whole equation) ?